OpenStudy (anonymous):

A parking lot charges $3 to park a car for the first hour and $2 per hour after that. If you use more than one parking space, the second and each subsequent car will be charged 75% of what you pay to park just one car. If you park 3 cars for t hours, which function gives the total parking charge?

OpenStudy (taylorterminate):

(3 + 2(t − 1))= price of first car for t hours 0.75(3 + 2(t − 1)) = price of second car with 75% off 0.75 0.75(3 + 2(t − 1))= price of third car, take first eq, take 75% off giving you the price of the second car but it said take ANOTHER 75% so you multiply by .75 once again going you the price of the third car for t hours .

OpenStudy (anonymous):

i believe there is a formula for compound interest, hang on 

OpenStudy (taylorterminate):

f(t) = (3 + 2(t − 1)) + 0.75(3 + 2(t − 1)) + 0.75 0.75(3 + 2(t − 1))